How much 25% hcl and 60% hcl will it take to make 200 ml of 45% hcl?
Answer by jorel1380(3719)

200(.45)=90 ml hcl
x+y=200
.25x+.60y=90
25x+60y=9000
25x+25y=5000
35y=4000
y=114.29
x=85.71
You need 85.71 ml of 25% hcl and 114.29 ml of 60% hcl to make 200 ml of 45% hcl..